Show an evaluation of sin 120 degrees using the trigonometric addition formulas.
Show an evaluation of sin 120 degrees using the double-angle formulas.

Answer by lwsshak3(11628)

Show an evaluation of sin 120 degrees using the trigonometric addition formulas.
Show an evaluation of sin 120 degrees using the double-angle formulas...
..
Addition formula:
sin(x+y)=sinxcosy+cosxsiny
sin120=sin(60+60)
=sin60cos60+cos60sin60
=(sqrt(3)/2)(1/2+(1/2)(sqrt(3)/2)
=(sqrt(3)/2)
..
Double angle formula:
Sin2x=2sinxcosx
Sin120=2sin60cos60
=2(sqrt3/2)(1/2)
=(sqrt3/2)